Target Profit

Scrushy Company sells a product for $150 per unit. The variable cost is $110 per unit, and fixed costs are $200,000.

Determine (a) the break-even point in sales units and (b) the break-even point in sales units if the company desires a target profit of $50,000.

a. Break-even point in sales units units
b. Break-even point in sales units if the company desires a target profit of $50,000
Answer #1

a) Break even point units = Fixed cost/Contribution margin per unit = 200000/(150-110) = 5000 Units

b) Break even point units if target profit is = (200000+50000)/40 = 6250 Units